OpenStudy (candy13106):

How many sides does a regular polygon have if the measure of an interior angle is 160 degree? Justify ur answer

OpenStudy (anonymous):

sum = 180(n-2) where n is # of sides so set up like 160 = 180 (n-2) solve for n

OpenStudy (candy13106):

thanks

OpenStudy (candy13106):

so wats the answer

OpenStudy (anonymous):

what did you get..when you solved for N?

OpenStudy (anonymous):

I am sorry, I read the question wrong. You were asked if "an" interior angle is 160... so use 160 = (n-2)180/n 160n = 180n - 360 360 = 20n 18 = n

OpenStudy (candy13106):

its fine but thx :]